A thermometer is removed from a room where the temperature is 70° F and is taken outside, where the air temperature is 40° F. After one-half minute the thermometer reads 60° F. What is the reading of the thermometer (in degrees Ferinheight) at t = 1 min?  (Round your answer to two decimal places.)

How long (in minutes) will it take for the thermometer to reach 45° F? (Round your answer to two decimal places.)
Answer #1

* Given that :- The toom temperature ... cor temperature Mhe. of = 40° F The Temparative after 142 min = 60°F 80 tel Min tempKo 10115 8s (/3) R = 283 On 93) That is from of T(t) = 40+20 203 16 (013] : 40+0 elen 1932)+ + 40+30 23ht Tht) = 40+20 (2)]@t